OpenStudy (anonymous):

integral x^3/(x^3 +1)

OpenStudy (anonymous):

\(\frac{x^3}{x^3+1}=1-\frac{1}{x^3+1}=1-\frac{1}{(x+1)(x^2-x+1)}\). Then you should use the partial fraction expansion to expand the second term.